# thanks: HJK and EGM are first co-authors of this manuscriptthanks: HJK and EGM are first co-authors of this manuscript

### v1

@@ I Introduction
In this communication, we highlight the potential and need for coherent lineshape analysis in $\pi$-conjugated polymers. We focus on a prototypical polymer poly(2,5-bis(3-hexadecylthiophene-2-yl)- thieno[3,2-b]thiophene) (PBTTT, Fig. 1a), which has a hairy-rod structure due to the fused rings in its backbone and long sidechain substitutions [31, 32, 33]. These two effects can induce liquid-crystalline-like behavior in PBTTT, producing a solid-state structure that often is highly structurally dynamic [34, 35, 36]. We first analyze the absorption and emission spectra of PBTTT using the framework of the weakly-coupled hybrid HJ-aggregate model. We then attempt to interpret the rephasing multidimensional coherent spectrum and its evolution with interpulse delay using the same framework. These 2D spectra are able to distinguish between homogeneous and inhomogeneous broadening, potentially revealing additional features hidden under the broad linear linewidths.
@@ II.1 Analysis of linear lineshapes
We first consider the lineshapes of the PBTTT linear absorption and photoluminescence (PL) spectra because they contain intricate information on the excitonic intra- and intermolecular coupling and energetic disorder [3]. Fig. 1b displays the linear absorption and photoluminescence spectra of a PBTTT film measured at 5 K. These spectra display lineshapes consistent with the hybrid HJ-aggregate model by Yamagata and Spano [2], in which the two-dimensional spatial coherence function of the vibronic Frenkel exciton depends on the interplay of intrachain and interchain resonance electronic Coulomb coupling [10].
@@ II.2 Coherent nonlinear spectroscopy
Figure 3: (a–d) Real and (e–h) imaginary part of the rephasing coherent 2D spectra of a PBTTT film, measured at 5 K, as a function of population waiting time $t_{\mathrm{pop}}$ labeled on the spectra.
@@ Acknowledgments
FT, HL, and CSA acknowledge funding from the Government of Canada (Canada Excellence Research Chair CERC-2022-00055). CSA acknowledges support from the Institut Courtois, Faculté des arts et des sciences, Université de Montréal (Chaire de Recherche de l’Institut Courtois) and from the Natural Science and Engineering Research Council of Canada (Discovery Grant RGPIN-2024-05893). HK acknowledges funding from the National Science Foundation Graduate Research Fellowship (DGE-2039655). Work at Georgia Tech (HK, NS, and CSA) was supported by the Center for Soft PhotoElectroChemical Systems, an Energy Frontier Research Center funded by the DOE, Office of Science, BES under Award # DE-SC0023411 (data analysis). ERB acknowledges funding from the National Science Foundation (CHE-2404788) and the Robert A. Welch Foundation (E-1337).
@@
The following author contribution abides to the CRediT contribution taxonomy. Natalie Stingelin, Carlos Silva-Acuña, Eric R. Bittner, Hao Li and Félix Thouin were responsible for the conceptualization of the project. Data curation and formal analysis were performed by Félix Thouin, Eliabeth Gutiérrez-Meza and Henry J. Kantrow. Carlos-Acuña, Natalie Stingelin, and Eric R. Bittner were responsible for funding acquisition. Henry J. Kantrow, Eliabeth Gutiérrez-Meza, Hongmo Li, Félix Thouin, and Hao Li were involved in the investigation. Félix Thouin and Carlos Silva-Acunã handled the project’s administration. The materials were synthesized by Martin Heeney and Qiao He; Hongmo Li was responsible for the polymer processing. Software and algorithms were developed by Félix Thouin, Henry J. Kantrow, and Eliabeth Gutiérrez-Meza. Carlos Silva-Acunã and Natalie Stingelin were responsible for the supervision of students involved in the project. Validation was performed by Henry J. Kantrow. Visualization of data was handled by Henry J. Kantrow and Eliabeth Gutiérrez-Meza. Henry Kantrow, Eliabeth Gutiérrez-Meza, Félix Thouin, Hao Li and Carlos Silva were involved in writing the original draft. All authors were involved in review and editing of the manuscript.
@@
Data and code used to produce the figures in this article are publicly available on the Borealis repository under DOI XXXXXXXXX.
@@
PBTTT-C16 was synthesized as previously reported ($M_{n}=38$ kg mol-1, $D=1.78$[31] and dissolved at 10 mg mL-1 in 1,2-dichlorobenzene (o-DCB, Sigma-Aldrich) at 85C for 1 hour. Films were drop-cast on sapphire substrates preheated to $30^{\circ}$ and allowed to dry on the heated stage.
@@
Henry J. Kantrow, Elizabeth Gutiérrez-Meza, Hongmo Li, Natalie Stingelin, Eric R. Bittner, Carlos Silva-Acuña

### v3

@@ I Introduction
In this communication, we highlight the potential and need for coherent lineshape analysis in $\pi$-conjugated polymers. We focus on a prototypical polymer poly(2,5-bis(3-hexadecylthiophene-2-yl)- thieno[3,2-b]thiophene) (PBTTT, Fig. 1a), which has a hairy-rod structure due to the fused rings in its backbone and long sidechain substitutions [31, 32, 33]. These two effects can induce liquid-crystalline-like behavior in PBTTT, producing a solid-state structure that often is highly structurally dynamic [34, 35, 36, 37]. We first analyze the absorption and emission spectra of PBTTT using the framework of the weakly-coupled hybrid HJ-aggregate model. We then attempt to interpret the rephasing multidimensional coherent spectrum and its evolution with interpulse delay using the same framework. These 2D spectra are able to distinguish between homogeneous and inhomogeneous broadening, potentially revealing additional features hidden under the broad linear linewidths.
@@ II.1 Analysis of linear lineshapes
We first consider the lineshapes of the PBTTT linear absorption and photoluminescence (PL) spectra because they contain intricate information excitonic intra- and intermolecular coupling and energetic disorder [3]. Fig. 1 displays the linear absorption and photoluminescence spectra of a PBTTT film measured at 5 K. These spectra display lineshapes consistent with the hybrid HJ-aggregate model by Yamagata and Spano [2], in which the two-dimensional spatial coherence function of the vibronic Frenkel exciton depends on the interplay of intrachain and interchain resonance electronic Coulomb coupling [10].
@@ II.2 Coherent nonlinear spectroscopy
Figure 3: (a-d) Real and (e-h) imaginary part of the rephasing coherent 2D spectra of a PBTTT film, measured at 5 K, as a function of population waiting time $t_{\mathrm{pop}}$ labeled on the spectra.
@@ Acknowledgments
FT, HL, and CSA acknowledge funding from the Government of Canada (Canada Excellence Research Chair CERC-2022-00055). CSA acknowledges support from the Institut Courtois, Faculté des arts et des sciences, Université de Montréal (Chaire de Recherche de l’Institut Courtois) and from the Natural Science and Engineering Research Council of Canada (Discovery Grant RGPIN-2024-05893). HK acknowledges funding from the National Science Foundation Graduate Research Fellowship (DGE-2039655). Work at Georgia Tech (HK, NS, and CSA) was supported by the Center for Soft PhotoElectroChemical Systems, an Energy Frontier Research Center funded by the DOE, Office of Science, BES under Award # DE-SC0023411 (data analysis). ERB acknowledges funding from the National Science Foundation (CHE-2404788) and the Robert A. Welch Foundation (E-1337). MH acknowledges funding from the EPSRC (EP/T028513/1).
@@
The following author contribution abides to the CRediT contribution taxonomy. Natalie Stingelin, Carlos Silva-Acuña, Eric R. Bittner, Hao Li and Félix Thouin were responsible for the conceptualization of the project. Data curation and formal analysis were performed by Félix Thouin, Elizabeth Gutiérrez-Meza and Henry J. Kantrow. Carlos-Acuña, Natalie Stingelin, and Eric R. Bittner were responsible for funding acquisition. Henry J. Kantrow, Elizabeth Gutiérrez-Meza, Hongmo Li, Félix Thouin, and Hao Li were involved in the investigation. Félix Thouin and Carlos Silva-Acunã handled the project’s administration. The materials were synthesized by Martin Heeney and Qiao He; Hongmo Li was responsible for the polymer processing. Software and algorithms were developed by Félix Thouin, Henry J. Kantrow, and Elizabeth Gutiérrez-Meza. Carlos Silva-Acunã and Natalie Stingelin were responsible for the supervision of students involved in the project. Validation was performed by Henry J. Kantrow. Visualization of data was handled by Henry J. Kantrow and Elizabeth Gutiérrez-Meza. Henry Kantrow, Elizabeth Gutiérrez-Meza, Félix Thouin, Hao Li and Carlos Silva were involved in writing the original draft. All authors were involved in review and editing of the manuscript.
@@
Data and code used to produce the figures in this article are publicly available on the Borealis repository [53].
@@
PBTTT-C16 was synthesized as previously reported ($M_{n}=38$ kg mol-1, Đ $=1.78$[31] and dissolved at 10 mg mL-1 in 1,2-dichlorobenzene (o-DCB, Sigma-Aldrich) at 85C for 1 hour. Films were drop-cast on sapphire substrates preheated to $30^{\circ}$ and allowed to dry on the heated stage.
@@
Henry J. Kantrow, Elizabeth Gutiérrez-Meza, Hongmo Li, Qiao He, Martin Heeney, Natalie Stingelin, Eric R. Bittner, Carlos Silva-Acuña, Hao Li, and Félix Thouin